OpenStudy (anonymous):

Please help.What is the volume of the regular pyramid below? A. 2700 units3 B. 8100 units3 C. 270 units3 D. 810 units3

OpenStudy (anonymous):

umm i dont see pyramid

OpenStudy (anonymous):



OpenStudy (jdoe0001):

area of a square pyramid = \(\bf \cfrac{1}{3} \times \text{squarish base area} \times\text{height}\)

OpenStudy (anonymous):

i got 2699.99~ and rounded to 2700 uunits^3

OpenStudy (jdoe0001):

what's the "squarish base" Area?

OpenStudy (anonymous):

100

OpenStudy (jdoe0001):

and the height?

OpenStudy (anonymous):

81

OpenStudy (jdoe0001):

\(\bf \cfrac{1}{3} \times 100 \times 81\)

OpenStudy (anonymous):

wich gave me 2699.9 and rounded to 2700

OpenStudy (jdoe0001):

2699.9?

OpenStudy (anonymous):

2699.9~

OpenStudy (anonymous):

well i guess my phone calculator wass close? but thanks i got it right (:

OpenStudy (jdoe0001):

yw